I need help please with these problems: Solve the absolute value equation and graph the solution set:
[2x-8]=0
[5x-1]<=14

Answer by stanbon(75887)

Solve the absolute value equation and graph the solution set:
[2x-8]=0
2x-8=0
x=4
----------------
[5x-1]<=14
-14<=5x-1<=14
Add 1 along the line to get:
-13 <=5x<=15
Divide by 5 along the line to get:
-13/5 <= x <= 3
================
Cheers,
Stan H.